# A scalable adaptive deep Koopman predictive controller for real-time optimization of mixed traffic flow

### v1

@@ III-A Key terms definition
III-B Problem description
Assuming that in high-density mixed traffic flow, HDVs not engaging in lane-changing behavior are primarily influenced by their preceding vehicles. Their driving behavior is primarily shaped by the current velocity of the preceding vehicle, their own current velocity, and the headway. For the purpose of research, time is discretized into infinitesimally small segments, and the aforementioned process can be described by the following equations:
@@ III-B Problem description
where , , , respectively represent the velocity difference, acceleration, and vehicle length of vehicle ; , , respectively represent the length of historical trajectories, the mapping relationship between trajectories and driving characteristics, and the driving characteristics extracted from information containing trajectory samples.
@@
III-C Koopman operator theory for state prediction of HDVs
III-C1 Koopman operator theory
The Koopman operator theory initially provides an alternative linear dynamic description for the evolution of uncontrollable systems [39]. With slight modifications, the Koopman operator can be applied to controlled systems [45]. Therefore, the evolution of system modeled by Eq. (2) can be expressed by a linear Koopman operator in an infinite-dimensional space. Let represents the state of vehicle , denotes all the velocities in the velocity space , the Koopman operator on System corresponding to Eq.(2) with the extended state is defined as follows:
@@ III-C1 Koopman operator theory
III-C2 Extended dynamic mode decomposition
The extended dynamic mode decomposition (EDMD) is a data-driven approach of finding the finite-dimensional approximation of the Koopman operator [46]. EDMD employs various basis functions, such as Radial Basis Functions (RBF) with different kernel centers and widths, to represent observable functions. It utilizes least squares regression to calculate . For forced dynamics model in Eq. (2), a special way for selecting basis functions is defined in Eq. (5) to obtain :
@@ III-D Model architecture
The model architecture is depicted in Fig.2. For trajectory context inputs, the model incorporates a driving characteristic semantic extraction block (highlighted by the deep green dashed box in Fig.2). To handle the current vehicle state input, a multi-layer perceptron-based encoder-lifting function approximation is employed. The fusion gate mechanism integrates the encoding of driving characteristics with the explicit state encoding of the lifting space, yielding the state of the observable high-dimensional Koopman space approximation. By incorporating the future velocity of the preceding vehicle into the network, the model adaptively learns the Koopman operator approximation and achieves multi-step state predictions in the high-dimensional space. This module is denoted as the Koopman evolution block, highlighted by the pink dashed box in Fig.2. Additionally, a linear decoder, serving as an approximation for Koopman modes, is utilized to transform observations from the high-dimensional space to obtain predicted states in the original space.
III-D1 Driving characteristic semantic extraction block
@@
Input Embedding and Temporal Encoding (ITE) Module The input embedding layer serves to convert the trajectory context into a high-dimensional dense representation, allowing the model to comprehensively learn the trajectory features and analyze correlation between trajectory samples. The embedding is achieved through a fully connected layer with a ReLU activation function. Given that the embedding operation is conducted for each trajectory sample, a relative time encoding method is introduced to enable subsequent modules to recognize the temporal information of trajectory context[47]. The expression is as follows:
where , represent the encoding feature dimension index and time step, respectively; represents the dimensionality of the encoding of each module in the AdapKoopnet, without special declaration. Subsequently, the trajectory context input embedding and temporal encoding are added together to form the output of this module. The expression is as follows:
@@
where , , , represents learnable weights; denotes the concatenate operation. stands for Layer Normalization, a technique that normalizes the trajectory encoding along the feature dimension to mitigate the impact of internal covariate shift [48]. Unlike Batch Normalization (BN), Layer Normalization is more flexible as it is not influenced by the size of the data batch.
After achieving feature exchange within the trajectory context through the attention layer, the feedforward layer is employed for the nonlinear transformation of trajectory encoding. This aims to capture the nonlinear relationships within the trajectory context, facilitating the model in learning higher-level abstract representations. The feedforward layer consists of two linear transformations and an activation function. Initially, undergoes a fully connected layer, followed by the application of the ReLU activation function, and finally passes through another fully connected layer. After goes through the feedforward layer, the output of the DTI module is obtained, calculated using the following formula:
@@
The process begins by defining learnable feature weights corresponding to each scenario. These weights dynamically learn the importance of each feature in the trajectory context encoding under predefined scenarios, using an extensive training data. The features are then aggregated based on the learned importance, obtaining the relevance of each trajectory sample to driving characteristic extraction in a specific driving scenario. Then, the relevance for each scenario is aggregated based on the probabilities predicted by the DSR module or according to user-defined scenarios. After applying the softmax activation function, the correlation between the trajectory context encoding and driving features is obtained. Based on this correlation, the trajectory context is aggregated, ultimately revealing the driving features hidden within the trajectory context. The computational formula is as follows:
@@
III-D2 Explicit state encoder and fusion gate mechanism
The explicit state encoder is tasked with encoding the explicit state of the vehicle from the original space to an observable high-dimensional space, facilitating the action of the Koopman lifting function on the current vehicle state. This encoder initially embeds the vehicle state into a high-dimensional representation through a linear layer with a ReLU activation function. Subsequently, a Multi-Layer Perceptron (MLP) with tanh activation function is utilized to perform multiple linear transformations on the embedded representation, completing the task of encoding the explicit state of the vehicle. Next, the explicit state encoding and the latent driving characteristics of the vehicle need to be fused to form the Koopman state approximation for the current time step. The Gated Linear Unit (GLU) is introduced to accomplish this fusion task. And the computational formula is as follows:
where ; , represents the weights of explicit state encoder and fusion gate mechanism. , are the explicit state encoding and Koopman state approximation of the vehicle at time step .
III-D3 Koopman evolution block and decoder
@@
III-E Loss function
Loss of AdapKoopnet is composed of reconstruction error, prediction error, and linear evolution error [49]. Specifically, the reconstruction error represents the difference between the reconstructed state obtained by embedding the current explicit state of the vehicle into a high-dimensional space and reconstructing it through the decoder, and the original state. To achieve accurate reconstruction of the original state, the reconstruction error for the entire prediction horizon is included as part of the loss function, expressed as follows:
@@ III-E Loss function
where , , are the weights corresponding to the three parts of the loss, and they are hyperparameters. To avoid manually selecting hyperparameters, Dynamic Weight Aver age (DWA) are introduced [50]. These weights are adjusted based on the losses from the previous epoch in the dynamic loss function.
@@
IV AdapKoopPC: Predictive control framework of CAVs for real-time optimizing mixed traffic flow based on Adakoopnet
IV-A Problem description and system state representation
@@
Consider a local traffic flow within a high-density mixed traffic flow, as illustrated in Fig. 4 consisting of HDVs and CAVs driving in the same lane. The leading vehicle is a HDV numbered as . For convenience, CAVs are sequentially numbered as . The HDVs located between CAV and are sequentially numbered as . The equivalence relationships exist: . Assuming that all HDV states in this local traffic flow can be collected through the onboard sensors of CAVs or a V2X (Vehicle-to-Everything) system. The control problem of the local traffic flow is defined as achieving the following objectives through the rolling optimization of control inputs for CAVs: (1) Minimizing the velocity difference between any two adjacent vehicles in the local traffic flow, while maintaining reasonable headway to alleviate traffic oscillations transmitted by the leading vehicle. (2) Ensuring that the velocities of vehicles in the local traffic flow are close to that of the leading vehicle, aiming to mitigate oscillations while maintaining traffic efficiency.
@@
For the convenience of constraint formulation, the local traffic flow state is defined as follows:
@@
where encompasses the headways, velocities of all vehicles in the local traffic flow, as well as the velocity differences of HDVs and accelerations of CAVs; is the position of the first CAV in the local traffic flow. The control input of the local traffic flow is defined as the jerk of CAVs, i.e., the rate of change of acceleration, denoted as:
@@
Remark 3: In large-scale high-density mixed traffic flow scenarios, the lane-changing frequency is low, areas without lane changes occur can be divided into multiple local traffic flow as described above. The local traffic flow lacks a strict organizational structure, requiring cooperation among a few neighboring CAVs to establish. Unlike most existing research, the achievement of control objectives of the local traffic flow does not depend on the so-called equilibrium velocity and equilibrium headway. Minimizing the velocity difference between adjacent vehicles is more readily accepted by HDV drivers, aligns with their own driving objectives in most cases. For scenarios with low penetration rates of V2X systems, simulation experiments relying solely on the onboard sensors of CAVs are conducted and tested in Section 6.
IV-B State predictive model of mixed traffic flow
@@
To enable the rapid convergence of the local traffic flow to the desired state while satisfying the constraints on local traffic flow states and control inputs, a predictive control framework is adopted. The framework assumes a set of feasible control input sequences within a specified region known as the control horizon . Based on the predictive model, a series of predictive states within the predictive horizon is obtained. By minimizing the error between the reference state and the predictive state, the optimal control input that satisfies constraints is determined. The current step control input is then applied to the CAVs, and this iterative process ensures that the control input at each step for the CAVs is at least a suboptimal solution. It is evident that in this framework, the accuracy of the local traffic flow state predictive model is crucial. Firstly, the CAVs in the local traffic flow are controllable, and their predictive state is obtained through the following vehicle kinematics equations:
@@
where represent the local traffic flow matrix and state matrix of CAVs, respectively. In this local traffic flow, there is always at least one CAV in front of any HDV. Therefore, AdapKoopnet can be cascaded to predict the states of all vehicles under specific control inputs of CAVs. This forms the basis for resolving traffic flow stop-and-go waves. The predictive process for all HDVs is illustrated in Fig. 5. Prior to the predictive process, it is necessary to obtain the high-dimensional representation of the current state of the vehicles using the encoding block of AdapKoopnet.
@@
However, the predictive process seems to be cumbersome, particularly when local traffic flow involves numerous vehicles. Therefore, the state predictive model needs to be integrated before initiating prediction to perform parallel prediction of all vehicle states within the local traffic flow. For any HDV in local traffic flow, the following relationship exists:
@@
where is the observation matrix, which is the linear decoder in AdapKoopnet. For the HDVs, when the preceding vehicle is also an HDV, the following equations are further obtained:
@@
where represents the row in the observation matrix used for observing velocity. Combining Eqs. (22)-(26), the predictive model for the local traffic flow is obtained:
@@
IV-C AdapKoopPC for mixed traffic flow
IV-C1 Constrained optimization problem of mixed traffic flow
Assuming the predictive horizon is the same as the control horizon and denoting it as , at time , the optimization objective for the mixed traffic flow defined in Section 4.1 can be formulated as minimizing the following cost function over the horizon to find an optimal set of control inputs :
where is the predictive state of the local traffic flow after time steps at time ; represents the reference state, it provides the reference values for the velocity of CAVs in the local traffic flow, namely the average velocities of their respective preceding vehicles over the time interval to ; and are diagonal matrices, representing the penalty weights for state and control input, respectively, where contains penalties for velocities of CAVs and velocity differences of HDVs, defined as follows:
Additionally, the local traffic flow states and control inputs must satisfy the following constraints:
These constraints respectively imply that the headways of all vehicles in the local traffic flow should remain within a reasonable range, velocities should stay within velocity limits, and the acceleration and control inputs of CAVs should adhere to vehicle dynamics constraints.
IV-C2 Standard quadratic programming of AdapKoopPC
To facilitate solving, the constrained optimization problem of mixed traffic flow can be reformulated as a standard quadratic programming problem. Building on the linear model described in Eq. (25), the predictive states among the future steps can be expressed as:
where ; ; ; .
V Experiment Part I: AdapKoopnet for car following trajectory prediction of HDVs
V-A Dataset description and evaluation metrics
The performance of AdapKoopnet model is evaluated on the HighD dataset, which is a large-scale naturalistic vehicle trajectory dataset collected from German highways. The car-following trajectory dataset is extracted in accordance with the criteria presented in [28], includes 18396432 records. And the root mean square error (RMSE) is selected as the evaluation metrics:
V-B Baseline models and experiment settings

### v3

@@ 3.1 Key terms definition
3.2 Problem description
Assuming that in high-density mixed traffic flow, HDVs not engaging in lane-changing behavior are primarily influenced by their preceding vehicles. Their driving behavior is primarily shaped by the current velocity of the preceding vehicle, their own current velocity, and the headway. For the purpose of research, time is discretized into infinitesimally small segments, and the aforementioned process can be described by the following equations(Saifuzzaman and Zheng, 2014):
@@ 3.2 Problem description
where , respectively represent the velocity difference, acceleration, and vehicle length of vehicle ; , , respectively represent the length of historical trajectories, the mapping relationship between trajectories and driving characteristics, and the driving characteristics extracted from information containing trajectory samples.
@@
3.3 Koopman operator theory for state prediction of HDVs
3.3.1 Koopman operator theory
The Koopman operator theory initially provides an alternative linear dynamic description for the evolution of uncontrollable systems (Koopman, 1931). With slight modifications, the Koopman operator can be applied to controlled systems (Proctor et al., 2018). Therefore, the evolution of system modeled by Eq. (2) can be expressed by a linear Koopman operator in an infinite-dimensional space. Let represents the state of vehicle , denotes all the velocities in the velocity space , the Koopman operator on System corresponding to Eq.(2) with the extended state is defined as follows:
@@ 3.3.1 Koopman operator theory
3.3.2 Extended dynamic mode decomposition
The extended dynamic mode decomposition (EDMD) is a data-driven approach of finding the finite-dimensional approximation of the Koopman operator (Williams et al., 2015). EDMD employs various basis functions, such as Radial Basis Functions (RBF) with different kernel centers and widths, to represent observable functions. It utilizes least squares regression to calculate . For forced dynamics model in Eq. (2), a special way for selecting basis functions is defined in Eq. (5) to obtain :
@@ 3.4 Model architecture
The model architecture is depicted in Fig.2. For trajectory context inputs, the model incorporates a driving characteristic semantic extraction block (highlighted by the deep green dashed box in Fig.2). To handle the current vehicle state input, a multi-layer perceptron-based encoder-lifting function approximation is employed. The fusion gate mechanism integrates the encoding of driving characteristics with the explicit state encoding of the lifting space, yielding the state of the observable high-dimensional Koopman space approximation. By incorporating the future velocity of the preceding vehicle into the network, the model adaptively learns the Koopman operator approximation and achieves multi-step state predictions in the high-dimensional space. This module is denoted as the Koopman evolution block, highlighted by the pink dashed box in Fig.2). Additionally, a linear decoder, serving as an approximation for Koopman modes, is utilized to transform observations from the high-dimensional space to obtain predicted states in the original space.
3.4.1 Driving characteristic semantic extraction block
@@
Input Embedding and Temporal Encoding (ITE) Module The input embedding layer serves to convert the trajectory context into a high-dimensional dense representation, allowing the model to comprehensively learn the trajectory features and analyze correlation between trajectory samples. The embedding is achieved through a fully connected layer with a ReLU activation function. Given that the embedding operation is conducted for each trajectory sample, a relative time encoding method is introduced to enable subsequent modules to recognize the temporal information of trajectory context (Vaswani et al., 2017). The expression is as follows:
where , represent the encoding feature dimension index and time step, respectively; represents the dimensionality of the encoding of each module in the AdaptKoopnet, without special declaration. Subsequently, the trajectory context input embedding and temporal encoding are added together to form the output of this module. The expression is as follows:
@@
where , , , represents learnable weights; denotes the concatenate operation. stands for Layer Normalization, a technique that normalizes the trajectory encoding along the feature dimension to mitigate the impact of internal covariate shift (Ba et al., 2016). Unlike Batch Normalization (BN), Layer Normalization is more flexible as it is not influenced by the size of the data batch.
After achieving feature exchange within the trajectory context through the attention layer, the feedforward layer (Fine, 2006) is employed for the nonlinear transformation of trajectory encoding. This aims to capture the nonlinear relationships within the trajectory context, facilitating the model in learning higher-level abstract representations. The feedforward layer consists of two linear transformations and an activation function. Initially, undergoes a fully connected layer, followed by the application of the ReLU activation function, and finally passes through another fully connected layer. After goes through the feedforward layer, the output of the DTI module is obtained, calculated using the following formula:
@@
The process begins by defining learnable feature weights corresponding to each scenario. These weights dynamically learn the importance of each feature in the trajectory context encoding under predefined scenarios, using an extensive training data. The features are then aggregated based on the learned importance, obtaining the relevance of each trajectory sample to driving characteristic extraction in a specific driving scenario. Then, the relevance for each scenario is aggregated based on the probabilities predicted by the DSR module or according to user-defined scenarios. After applying the softmax activation function, the correlation between the trajectory context encoding and driving features is obtained. Based on this correlation, the trajectory context is aggregated, ultimately revealing the driving features hidden within the trajectory context. The computational formula for this module is as follows:
@@
3.4.2 Explicit state encoder and fusion gate mechanism
The explicit state encoder is tasked with encoding the explicit state of the vehicle from the original space to an observable high-dimensional space, facilitating the action of the Koopman lifting function on the current vehicle state. This encoder initially embeds the vehicle state into a high-dimensional representation through a linear layer with a ReLU activation function. Subsequently, a Multi-Layer Perceptron (MLP) with tanh activation function is utilized to perform multiple linear transformations on the embedded representation, completing the task of encoding the explicit state of the vehicle. Next, the explicit state encoding and the latent driving characteristics of the vehicle need to be fused to form the Koopman state approximation for the current time step. The Gated Linear Unit (GLU) is introduced to accomplish this fusion task. It employs a gate mechanism to adaptively filter redundant features and retain essential features (Dauphin et al., 2017). The computational formula for the above process is as follows:
where ; , represents the weights of explicit state encoder and fusion gate mechanism, respectively. , represent the explicit state encoding and Koopman state approximation of the vehicle at time step .
3.4.3 Koopman evolution block and decoder
@@
3.5 Loss function
Loss of AdapKoopnet is composed of reconstruction error, prediction error, and linear evolution error (Lusch et al., 2018). Specifically, the reconstruction error represents the difference between the reconstructed state obtained by embedding the current explicit state of the vehicle into a high-dimensional space and reconstructing it through the decoder, and the original state. To achieve accurate reconstruction of the original state, the reconstruction error for the entire prediction horizon is included as part of the loss function, expressed as follows:
@@ 3.5 Loss function
where , , are the weights corresponding to the three parts of the loss, and they are hyperparameters. To avoid manually selecting hyperparameters, Dynamic Weight Aver age (DWA) are introduced (Liu et al., 2019). These weights are adjusted based on the losses from the previous epoch in the dynamic loss function.
@@
4 AdapKoopPC: Predictive control framework of CAVs for real-time optimizing mixed traffic flow based on Adakoopnet
4.1 Problem description and system state representation
@@
Consider a local traffic system within a high-density mixed traffic flow, as illustrated in Fig. 4 consisting of HDVs and CAVs driving in the same lane. The leading vehicle is a HDV numbered as . For convenience, CAVs are sequentially numbered as . The HDVs located between CAV and are sequentially numbered as . The equivalence relationships exist: . Assuming that all HDV states in this system can be collected through the onboard sensors of CAVs or a V2X (Vehicle-to-Everything) system. The control problem of this system is defined as achieving the following objectives through the rolling optimization of control inputs for CAVs: (1) Minimizing the velocity difference between any two adjacent vehicles in the system, while maintaining reasonable headway to alleviate traffic oscillations transmitted by the leading vehicle. (2) Ensuring that the velocities of vehicles in the system are close to that of the leading vehicle, aiming to mitigate oscillations while maintaining traffic efficiency.
@@
Based on the above two objectives and for the convenience of constraint formulation, the system state is defined as follows:
@@
where encompasses the headways, velocities of all vehicles in the system, as well as the velocity differences of HDVs and accelerations of CAVs; is the position of the first CAV in the system. The control input of the system is defined as the jerk of CAVs, i.e., the rate of change of acceleration, denoted as:
@@
Remark 3: In large-scale high-density mixed traffic flow scenarios, the lane-changing frequency is low, areas without lane changes occur can be divided into multiple systems as described above. The system lacks a strict organizational structure, requiring cooperation among a few neighboring CAVs to establish. Unlike most existing research, the achievement of control objectives of the system does not depend on the so-called equilibrium velocity and equilibrium headway. Minimizing the velocity difference between adjacent vehicles is more readily accepted by HDV drivers, aligns with their own driving objectives in most cases. For scenarios with low penetration rates of V2X systems, simulation experiments relying solely on the onboard sensors of CAVs are conducted and tested in Section 6.
4.2 State predictive model of mixed traffic system
@@
To enable the rapid convergence of the system to the desired state while satisfying the constraints on system states and control inputs, a predictive control framework is adopted. The framework assumes a set of feasible control input sequences within a specified region known as the control horizon . Based on the predictive model, a series of predictive states within the predictive horizon is obtained. By minimizing the error between the reference state and the predictive state, the optimal control input that satisfies constraints is determined. The current step control input is then applied to the CAVs, and this iterative process ensures that the control input at each step for the CAVs is at least a suboptimal solution. It is evident that in this framework, the accuracy of the system state predictive model is crucial. Firstly, the CAVs in the system are controllable, and their predictive state is obtained through the following vehicle kinematics equations:
@@
where represent the system matrix and state matrix of CAVs, respectively. In this system, there is always at least one CAV in front of any HDV. Therefore, AdapKoopnet can be cascaded to predict the states of all vehicles in the system under specific control inputs of CAVs. This forms the basis for resolving traffic flow stop-and-go waves in the framework. The predictive process for all HDVs is illustrated in Fig. 5. Prior to the predictive process, it is necessary to obtain the high-dimensional representation of the current state of the vehicles using the encoding block of AdapKoopnet.
@@
However, the predictive process seems to be cumbersome, particularly when system involves numerous vehicles. Therefore, the state predictive model needs to be integrated before initiating prediction to perform parallel prediction of all vehicle states within the system. For any HDV in system, the following relationship exists:
@@
where represents the observation matrix, which is the linear decoder in AdaptKoopnet. For the HDVs, where the preceding vehicle is also an HDV, the following equations are further obtained:
@@
where represents the row in the observation matrix used for observing velocity. Combining Eqs. (22)-(26), the predictive model for the system is obtained:
@@
4.3 AdapKoopPC for mixed traffic system
4.3.1 Constrained optimization problem of mixed traffic system
Assuming the predictive horizon is the same as the control horizon and denoting it as , at time , the optimization objective for the mixed traffic system defined in Section 4.1 can be formulated as minimizing the following cost function over the horizon to find an optimal set of control inputs :
where is the predictive state of the system after time steps at time ; represents the reference state, it provides the reference values for the velocity of CAVs in the system, namely the average velocities of their respective preceding vehicles over the time interval to ; and are diagonal matrices, representing the penalty weights for state and control input, respectively, where contains penalties for velocities of CAVs and velocity differences of HDVs, defined as follows:
Additionally, the system states and control inputs must satisfy the following constraints:
These constraints respectively imply that the headways of all vehicles in the system should remain within a reasonable range, velocities should stay within velocity limits, and the acceleration and control inputs of CAVs should adhere to vehicle dynamics constraints.
4.3.2 Standard quadratic programming of AdapKoopPC
To facilitate solving, the constrained optimization problem of mixed traffic system can be reformulated as a standard quadratic programming problem. Building on the linear model described in Eq.(25), the predictive states among the future steps can be expressed as:
where ; ; .
Let , , , where is Kronecker product. And by discarding the terms without , the cost function Eq.(29) has been rewritten as follows:
where ; . Optimal control input sequence can be obtained by solving the constrained optimization problem (33) subject to Eqs. (31) and (32).
5 Experiment Part I: AdapKoopnet for car following trajectory prediction of HDVs
5.1 Dataset description and evaluation metrics
The performance of AdapKoopnet model is evaluated on the HighD dataset, which is a large-scale naturalistic vehicle trajectory dataset collected from German highways. The dataset comprises 11.5 hours of measurements and covers 110,000 vehicles, with a total measured vehicle distance of 45,000 km, capturing trajectories of both passenger cars and trucks. The car-following trajectory dataset is extracted in accordance with the criteria presented in (Mo et al., 2021), includes eleven fields such as id, precedingId, xVelocity, xAcceleration, and precedingXVelocity, totaling 18396432 records. And the root mean square error (RMSE) is selected as the evaluation metrics:
5.2 Baseline models and experiment settings